# Adaptive Prompt: Unlocking the Power of Visual Prompt Tuning

### v4

@@ Abstract
Visual Prompt Tuning (VPT) has recently emerged as a powerful method for adapting pre-trained vision models to downstream tasks. By introducing learnable prompt tokens as task-specific instructions, VPT effectively guides pre-trained transformer models with minimal overhead. Despite its empirical success, a comprehensive theoretical understanding of VPT remains an active area of research. Building on recent insights into the connection between mixture of experts and prompt-based approaches, we identify a key limitation in VPT: the restricted functional expressiveness in prompt formulation. To address this limitation, we propose Visual Adaptive Prompt Tuning (VAPT), a new generation of prompts that redefines prompts as adaptive functions of the input. Our theoretical analysis shows that this simple yet intuitive approach achieves optimal sample efficiency. Empirical results on VTAB-1K and FGVC further demonstrate VAPT’s effectiveness, with performance gains of 7.34% and 1.04% over fully fine-tuning baselines, respectively. Notably, VAPT also surpasses VPT by a substantial margin while using fewer parameters. These results highlight both the effectiveness and efficiency of our method and pave the way for future research to explore the potential of adaptive prompts.
@@ 1 Introduction
Foundational vision models have demonstrated remarkable success across a wide range of computer vision tasks [8, 55, 32], primarily by leveraging large-scale pre-training for robust generalization. As a result, fine-tuning these models for specific downstream tasks has become a widely adopted paradigm [26]. However, fully fine-tuning large foundational models can be computationally prohibitive, leading to growing interest in parameter-efficient fine-tuning (PEFT) techniques [2, 76, 25], which update only a small subset of parameters. Among these methods, Visual Prompt Tuning (VPT) [28] has emerged as a simple yet powerful approach that appends learnable prompt tokens to the input, serving as task-specific instructions to guide the pre-trained transformer model. Despite VPT’s empirical effectiveness, its theoretical underpinnings remain an active area of research [54, 51, 67].
@@
Recently, [35] uncovered a novel connection between attention mechanisms [64], prompt-based methods, and Mixture of Experts (MoE) models [27, 59], offering fresh insights into the design and optimization of prompt-based approaches. Specifically, they show that each attention head in a transformer can be equivalently viewed as a composition of multiple MoE models stacked together. This finding implies that, within a pre-trained transformer, each attention head encodes a collection of pre-trained MoE experts whose parameters are embedded in the model’s pre-trained weights. Under this framework, prompt tuning corresponds to fine-tuning these pre-trained MoE models by introducing new, learnable prompt experts. These prompt experts collaborate with the pre-trained experts to enable effective task adaptation. This connection paves the way for deeper theoretical investigations and the development of advanced strategies for prompt-based learning [36, 34].
@@
Building on this perspective, we identify a key limitation in the current formulation of VPT. Specifically, while pre-trained experts within attention heads are linear functions of the input features, the newly introduced prompt experts are constant functions that do not vary with the input. Since prompt experts must collaborate with the pre-trained ones to achieve optimal downstream performance, we hypothesize that this restricted functional expressiveness may inherently limit the efficacy of prompt tuning. This limitation naturally raises an important question about the potential of adaptive prompts, where prompts are modeled as adaptive functions of the input. Although increasing adaptivity may lead to concerns about parameter overhead, we propose a novel formulation that strikes a balance between flexibility and memory efficiency. We refer to this methodology as Visual Adaptive Prompt Tuning (VAPT).
Our design comprises two main components: token-wise projectors and a shared feature projector, which leverages global information from input features to generate adaptive prompt tokens. Theoretical analysis shows that our formulation achieves optimal sample efficiency for prompt estimation. Empirical results further substantiate this claim. For instance, on the Stanford Dogs dataset [31], using only 1% of the data, while VPT achieves a modest accuracy of 3.6%, VAPT demonstrates a remarkable improvement with an accuracy of 60.1%. Moreover, extensive experiments across multiple benchmarks consistently show that VAPT outperforms VPT while utilizing fewer parameters, highlighting the effectiveness of its simple yet elegant design. Notably, VAPT also achieves state-of-the-art performance compared to other PEFT methods across different pre-trained objectives [4, 20], indicating robustness both theoretically and empirically.
Contribution. The main contributions of this work are summarized as follows: 1. By leveraging the connection between mixture of experts and prompt-based techniques, we identify a key limitation in the formulation of VPT: the restricted expressiveness of prompts, which remain invariant regardless of the input. 2. To address this limitation and explore the potential of adaptive prompts, we propose VAPT, a novel formulation that balances efficiency and flexibility. 3. Through theoretical analysis, we demonstrate that our formulation achieves optimal sample efficiency for prompt estimation, establishing a solid foundation for its practical effectiveness. 4. Extensive experiments demonstrate that VAPT consistently outperforms VPT with fewer parameters, and achieves state-of-the-art results compared to other commonly used PEFT methods.
@@
Organization. The paper proceeds as follows. In Section 2 provides background on visual prompt tuning and mixture of experts. In Section 3, we discuss the connection between VPT and MoE, highlighting the limited expressiveness of the current VPT formulation. Section 4 describes in detail the design of our method to achieve superior performance with fewer parameters. We present our theoretical analysis in Section 5, followed by our empirical results in Section 6. Finally, we conclude the paper and suggest potential directions for future research in Section 7.
@@
Notation. For any $n\in\mathbb{N}$, let $[n]=\{1,2,\ldots,n\}$. For a set $S$, $|S|$ denotes its cardinality. Given a vector $u=(u_{1},u_{2},\ldots,u_{d})\in\mathbb{R}^{d}$ and $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, we set $u^{\alpha}=u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}\ldots u_{d}^{\alpha_{d}}$, $|u|=u_{1}+u_{2}+\ldots+u_{d}$ and $\alpha!=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$. Moreover, $\|u\|$ denotes the Euclidean norm of $u$. For two positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$ for sone constant $C>0$. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates $a_{n}/b_{n}$ is stochastically bounded.
Figure 1: Illustration of the forward propagation of VPT across ViT blocks. Red crosses denote output tokens that are discarded. In each block, the prompt tokens ${\bm{P}}^{(l)}$ are appended to the input $\tilde{{\bm{X}}}^{(l)}$. The output tokens at the prompt positions are then replaced by the prompt tokens ${\bm{P}}^{(l+1)}$ of the next block before being propagated forward.
@@ 2 Background
In this section, we first discuss Visual Prompt Tuning in Section 2.1 and then Mixture of Experts model in Section 2.2. Additional related work is provided in Appendix C.
@@ 2.1 Visual Prompt Tuning
Vision Transformer (ViT) [8] has been shown to be highly effective as a backbone architecture for various visual recognition tasks. Typically, a ViT consists of $L$ transformer blocks, each containing a multi-head self-attention (MSA) layer followed by a feed-forward network (FFN). For clarity, we focus on a single transformer block. Let $\tilde{{\bm{X}}}^{(l)}=\left[{\bm{x}}_{1}^{(l)},\dots,{\bm{x}}_{N}^{(l)}\right]^{\top}\in\mathbb{R}^{N\times d}$ be the sequence of input tokens in the $l$-th ViT block, where ${\bm{x}}_{i}^{(l)}\in\mathbb{R}^{d}$ denotes the $i$-th token, $N$ is the sequence length, and $d$ is the embedding dimension. The MSA layer processes ${\bm{X}}^{Q}={\bm{X}}^{K}={\bm{X}}^{V}=\tilde{{\bm{X}}}^{(l)}\in\mathbb{R}^{N\times d}$ as query, key, and value, respectively. Its output can then be written as:
@@
for $m\in[M]$, where $M$ is the number of attention heads, $W^{O}\in\mathbb{R}^{Md_{v}\times d}$, and $W_{m}^{Q}\in\mathbb{R}^{d\times d_{k}}$, $W_{m}^{K}\in\mathbb{R}^{d\times d_{k}}$, $W_{m}^{V}\in\mathbb{R}^{d\times d_{v}}$ are projection matrices with $d_{k}=d_{v}=\frac{d}{M}$.
Recently, Visual Prompt Tuning (VPT) [28] has emerged as a promising method for fine-tuning pre-trained vision models. VPT introduces prompt parameters, ${\bm{P}}^{(l)}=\left[\bm{p}_{1}^{(l)},\dots,\bm{p}_{N_{p}}^{(l)}\right]^{\top}\in\mathbb{R}^{N_{p}\times d}$, which are appended to the input token sequence of ViT block. Here, $N_{p}$ denotes the prompt length. The inclusion of these prompts modifies the MSA layer’s output, which can be expressed as follows:
@@
$\quad\ \ =\left[\tilde{{\bm{h}}}_{m,1},\dots,\tilde{{\bm{h}}}_{m,N+N_{p}}\right]^{\top}\in\mathbb{R}^{(N+N_{p})\times d_{v}},$ (2)
with $m\in[M]$. During training, only the prompt parameters ${\bm{P}}^{(l)}$ and the classification head are optimized, while the pre-trained ViT parameters, including $W_{m}^{Q}$, $W_{m}^{K}$, $W_{m}^{V}$ and $W_{O}$, remain frozen. After the MSA layer, each output token is independently processed by the pre-trained FFN, with no interaction between tokens. In VPT, the output tokens of each block, where prompts act as queries, are replaced with new trainable prompts ${\bm{P}}^{(l+1)}$ for the subsequent block, as illustrated in Figure 1. Consequently, in equation (2), there is no need to compute $\tilde{{\bm{h}}}_{m,N+1},\dots,\tilde{{\bm{h}}}_{m,N+N_{p}}$. Omitting these computations does not affect the output tokens of the current block, and if no new prompts are introduced in the next block, these output tokens can simply be discarded.
@@ 2.2 Mixture of Experts
Mixture of Experts (MoE) is a class of statistical machine learning frameworks that combines multiple models, known as experts, to construct more expressive and accurate predictions [27, 30]. An MoE model typically consists of $N^{\prime}$ expert functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$ for $i\in[N^{\prime}]$, along with a gating function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N^{\prime}}$ which assigns weights to the experts based on learned score functions $s_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. For a given input ${\bm{h}}\in\mathbb{R}^{d}$, the MoE model generates its output as:
$\hat{\mathbf{y}}=\sum_{j=1}^{N^{\prime}}G({\bm{h}})_{j}\cdot f_{j}({\bm{h}})=\sum_{j=1}^{N^{\prime}}\frac{\exp\left(s_{j}({\bm{h}})\right)}{\sum_{\ell=1}^{N^{\prime}}\exp\left(s_{\ell}({\bm{h}})\right)}\cdot f_{j}({\bm{h}}),$
where $G({\bm{h}})=\mathrm{softmax}(s_{1}({\bm{h}}),\ldots,s_{N^{\prime}}({\bm{h}}))$. Recent studies have revealed a novel connection between the attention mechanism, prompt-based methods, and MoE frameworks [35, 36]. This connection presents promising opportunities for investigating prompt-based techniques through the lens of mixture of experts.
@@ 3 Motivation
Mixture of Experts meets Visual Prompt Tuning. We begin by discussing the connection between VPT and MoE. We adopt the same notation as in Section 2.1. From equation (2), let ${\bm{X}}^{(l)}=\left[{{\bm{x}}_{1}^{(l)}}^{\top},\dots,{{\bm{x}}_{N}^{(l)}}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ represent the concatenation of all $N$ input image tokens for the MSA layer. For notational simplicity, we omit the superscript $(l)$ in the derivations below. Recent studies [35] have demonstrated that each output vector $\tilde{{\bm{h}}}_{m,i}$ for $i\in[N]$ within the $m$-th attention head can be equivalently expressed as the output of an MoE model, where ${\bm{X}}$ serves as the input. Specifically, we define a set of experts as follows:
@@
for $i\in[N]$. Here, $E_{j}\in\mathbb{R}^{d\times Nd}$ is a standard basis selector such that $E_{j}{\bm{X}}={\bm{x}}_{j}$. Finally, the output of the $m$-th attention head for the $i$-th token can be expressed as follows:
@@
From these formulations, we can see that each attention head in ViT implicitly encodes multiple MoE models $\tilde{{\bm{h}}}_{m,i}$ for $i\in[N]$. Unlike conventional MoE layers [59], where each expert and its score function operate on a single token embedding ${\bm{x}}_{i}$, the expert networks and score functions here process the entire input sequence ${\bm{X}}$. Furthermore, the experts $f_{1},\dots,f_{N}$ and their associated score functions are pre-trained alongside the ViT, as their parameters are already embedded within the pre-trained model (see equation (3) and equation (5)), thereby serving as pre-trained experts. Meanwhile, the new prompt experts $f_{N+1},\dots,f_{N+N_{p}}$, whose learnable parameters are contained in prompts (see equation (4) and equation (6)), are introduced to efficiently adapt the model for downstream tasks. In other words, one can view prompt tuning as an efficient method of fine-tuning these implicitly encoded MoE models by adding new prompt experts. These newly introduced prompt experts effectively act as downstream-task experts, enabling the model to specialize in new tasks without retraining the entire network.
@@
Prompts remain invariant to the input. From equation (4), it follows that prompt experts $f_{N+1},\dots,f_{N+N_{p}}$, although formally functions of ${\bm{X}}$, are represented by fixed prompt vectors that remain constant regardless of the input. In contrast, pre-trained experts $f_{1},\dots,f_{N}$ are linear functions of ${\bm{X}}$ (see equation (3)), making them comparatively more expressive. We hypothesize that this limited flexibility may diminish the effectiveness of prompt tuning as a fine-tuning strategy. Supporting this view, prior work [54] demonstrates that prompt tuning can only add a bias term to the output of an attention block, thereby restricting the model’s representational capacity. This observation naturally leads to the following question: Can the performance of visual prompt tuning be improved by making prompts more flexible and adaptive to the input? To address this, we propose a novel prompt formulation to enhance their adaptability under varying inputs in Section 4.
@@
The challenge of staying efficient. Despite the limitations noted above, one of the main advantages of the current prompt design in VPT is its simplicity. As indicated by equation (4), each prompt expert requires only $d$ parameters, making it highly parameter-efficient. One might consider a naive linear design for a prompt expert $f_{N+j^{\prime}}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$, where $f_{N+j^{\prime}}({\bm{X}})=W_{j^{\prime}}^{\top}{\bm{X}}$. However, this approach would introduce up to $Nd\times d_{v}$ parameters, which is $Nd_{v}$ times larger than the bias-expert formulation and would significantly increase computational costs. Consequently, this highlights the challenge of balancing input adaptability with cost-effectiveness, thereby ensuring that the benefits of a parameter-efficient fine-tuning technique are preserved.

### v5

# On the Expressiveness of Visual Prompt Experts

@@ Abstract
Visual Prompt Tuning (VPT) has proven effective for parameter-efficient adaptation of pre-trained vision models to downstream tasks by inserting task-specific learnable prompt tokens. Despite its empirical success, a comprehensive theoretical understanding of VPT remains an active area of research. Building on the recently established connection between Mixture of Experts (MoE) and prompt-based methods, wherein each attention head can be conceptualized as a composition of multiple MoE models, we reinterpret VPT as the introduction of new prompt experts into these MoE structures. We identify a key limitation in existing VPT frameworks: the restricted functional expressiveness of prompt experts, which remain static and thus limited in their adaptability. To address this, we propose Visual Adaptive Prompt Tuning (VAPT), a novel method that endows prompt experts with enhanced expressiveness while preserving parameter efficiency. Empirical evaluations on VTAB-1K and FGVC demonstrate that VAPT achieves substantial performance improvements, surpassing fully fine-tuned baselines by 7.34% and 1.04%, respectively. Moreover, VAPT consistently outperforms VPT while requiring fewer additional parameters. Furthermore, our theoretical analysis indicates that VAPT achieves optimal sample efficiency. Collectively, these results underscore the theoretical grounding and empirical advantages of our approach.
@@ 1 Introduction
Foundational vision models, pre-trained on large-scale datasets dosovitskiy2020image ; radford2021learning ; kirillov2023segment , have demonstrated remarkable success and robust generalization across a wide range of computer vision tasks. As a result, fine-tuning these models for specific downstream tasks has become a widely adopted paradigm iofinova2022well . However, fully fine-tuning large foundational models can be computationally prohibitive, leading to growing interest in parameter-efficient fine-tuning (PEFT) techniques cai2020tinytl ; zhang2020side ; hu2021lora , which update only a small subset of parameters. Among these methods, Visual Prompt Tuning (VPT) jia2022visual has emerged as a simple yet powerful approach that appends learnable prompt tokens to the input, which serve as task-specific instructions to guide the pre-trained transformer model. Despite VPT’s empirical effectiveness, its theoretical underpinnings remain an active area of research petrov2023prompting ; oymak2023role ; wang2024universality .
@@
Recently, le2024mixture established a formal connection between attention mechanisms vaswani2017attention , prompt-based methods, and Mixture of Experts (MoE) models Jacob_Jordan-1991 ; Quoc-conf-2017 , yielding new insights into the design and optimization of prompting strategies. Their analysis demonstrates that each attention head in a transformer can be equivalently interpreted as a composition of multiple MoE models stacked together. Within this framework, VPT corresponds to fine-tuning these implicit, pre-trained MoE structures by introducing new, learnable prompt experts. These prompt experts collaborate with the pre-trained experts to facilitate effective task adaptation. This connection opens avenues for deeper theoretical investigations and the development of advanced strategies for prompt-based learning le2024revisiting ; le2024adaptive ; diep2025zero .
@@
Building on this MoE interpretation, we identify a key limitation in the standard formulation of VPT. While the pre-trained experts within attention heads are linear functions of the input features, the newly introduced prompt experts are modeled as constant, input-invariant vectors. Given that effective adaptation relies on collaboration between these prompt and pre-trained experts, we hypothesize that this functional disparity, specifically, the restricted expressiveness of static prompts, may constrain VPT’s adaptation efficacy. Furthermore, this design deviates from standard MoE practices, where experts are typically designed to be input-adaptive. This observation naturally raises an important question about the potential of adaptive prompts, wherein prompts are modeled as more expressive functions of the input. Although enhancing adaptivity might raise concerns about parameter overhead, we propose a novel formulation that injects input-dependent prompts while preserving the parameter efficiency characteristic of VPT. We term this approach Visual Adaptive Prompt Tuning (VAPT).
The VAPT design comprises two main components: token-wise projectors and a shared feature projector, which leverage global information from input features to generate adaptive prompt tokens. A key advantage of VAPT is its efficiency: it achieves enhanced expressiveness with minimal computational overhead, adding only 0.6% FLOPs relative to VPT and requiring fewer trainable parameters. Additionally, the token-wise projectors and channel-wise convolutions result in a structurally simple formulation of the prompt experts. This simplicity enables a rigorous theoretical analysis (see Section 5), a contribution largely absent in the prior prompt-tuning literature. Our analysis demonstrates that VAPT attains optimal sample efficiency for prompt estimation. Empirical results strongly corroborate these theoretical findings. For example, in a low-data regime on the Stanford Dogs dataset khosla2011novel (using only 1% of training data), VAPT achieves 60.1% accuracy, a stark contrast to VPT’s 3.6%. Moreover, on standard benchmarks VTAB-1K zhai2019large and FGVC jia2022visual , VAPT significantly improves performance over fully fine-tuned baselines by 7.34% and 1.04%, respectively. Crucially, VAPT consistently outperforms VPT across benchmarks despite utilizing fewer trainable parameters. These results underscore the theoretical soundness and empirical effectiveness of VAPT.
Contributions. 1. From MoE perspective, we identify a key limitation in the formulation of VPT: its prompt experts are input-invariant, thereby constraining their functional expressiveness. 2. We introduce VAPT, a novel formulation that injects input-adaptive prompt experts while preserving the parameter-efficiency of VPT. 3. VAPT’s simple formulation enables a theoretical analysis to validate its effectiveness, an aspect largely missing in prior work. Our theoretical analysis demonstrates that VAPT attains optimal sample efficiency for prompt estimation, providing a rigorous foundation for its practical value. 4. Extensive experiments show that VAPT consistently surpasses VPT with fewer trainable parameters, validating its effectiveness and efficiency both theoretically and empirically.
@@
Organization. Section 2 reviews visual prompt tuning and mixture of experts. In Section 3, we discuss the connection between VPT and MoE, highlighting the limited expressiveness of the standard VPT formulation. Section 4 details our proposed method, designed to achieve higher accuracy with greater parameter efficiency. We present our theoretical analysis in Section 5, followed by empirical results in Section 6. Finally, Section 7 concludes the paper and suggests potential future directions.
@@
Notation. For $n\in\mathbb{N}$, let $[n]=\{1,2,\ldots,n\}$. For a set $S$, $|S|$ denotes its cardinality. Given a vector $u=(u_{1},u_{2},\ldots,u_{d})\in\mathbb{R}^{d}$ and $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, define $u^{\alpha}=u_{1}^{\alpha_{1}}u_{2}^{\alpha_{2}}\ldots u_{d}^{\alpha_{d}}$, $|u|=u_{1}+u_{2}+\ldots+u_{d}$ and $\alpha!=\alpha_{1}!\alpha_{2}!\ldots\alpha_{d}!$. Let $\|u\|$ denote the Euclidean norm of $u$. For positive sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}$, we write $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for all $n\in\mathbb{N}$ for some $C>0$. The notation $a_{n}=\mathcal{O}_{P}(b_{n})$ indicates $a_{n}/b_{n}$ is stochastically bounded.
@@ 2 Background
In this section, we first review Visual Prompt Tuning in Section 2.1 and then the Mixture of Experts model in Section 2.2. Additional related work is presented in Appendix C.
@@ 2.1 Visual Prompt Tuning
Vision Transformer (ViT) dosovitskiy2020image has proven to be a powerful backbone architecture for visual recognition. A ViT contains $L$ blocks, each comprising a multi-head self-attention (MSA) layer followed by a feed-forward network (FFN). For clarity, we consider the $l$-th ViT block. Let $\tilde{{\bm{X}}}^{(l)}=\left[{\bm{x}}_{1}^{(l)},\dots,{\bm{x}}_{N}^{(l)}\right]^{\top}\in\mathbb{R}^{N\times d}$ be the input tokens, where $N$ is the sequence length, and $d$ is the embedding dimension. The MSA layer processes ${\bm{X}}^{Q}={\bm{X}}^{K}={\bm{X}}^{V}=\tilde{{\bm{X}}}^{(l)}\in\mathbb{R}^{N\times d}$ as queries, keys, and values, producing:
@@
for $m\in[M]$, where $M$ is the number of attention heads, $W_{m}^{Q}\in\mathbb{R}^{d\times d_{k}}$, $W_{m}^{K}\in\mathbb{R}^{d\times d_{k}}$, $W_{m}^{V}\in\mathbb{R}^{d\times d_{v}}$, and $W^{O}\in\mathbb{R}^{Md_{v}\times d}$ are projection matrices with $d_{k}=d_{v}=\frac{d}{M}$.
Visual Prompt Tuning (VPT) jia2022visual adapts a pre-trained ViT by appending $N_{p}$ learnable prompt parameters ${\bm{P}}^{(l)}=\left[\bm{p}_{1}^{(l)},\dots,\bm{p}_{N_{p}}^{(l)}\right]^{\top}\in\mathbb{R}^{N_{p}\times d}$, (where $N_{p}$ is the prompt length), to the input of each ViT block, thereby modifying the MSA layer’s output as follows:
@@
$\quad\ \ =\left[\tilde{{\bm{h}}}_{m,1},\dots,\tilde{{\bm{h}}}_{m,N+N_{p}}\right]^{\top}\in\mathbb{R}^{(N+N_{p})\times d_{v}}.$ (2)
During training, only the prompts ${\bm{P}}^{(l)}$ and the classification head are updated, while all ViT weights, including $W_{m}^{Q}$, $W_{m}^{K}$, $W_{m}^{V}$, and $W_{O}$, remain frozen. Within the VPT framework, MSA output tokens corresponding to the input prompt locations are discarded and replaced by the prompts of the next layer ${\bm{P}}^{(l+1)}$, before these tokens enter the subsequent block. Consequently, in Equation (2), $\tilde{{\bm{h}}}_{m,N+1},\dots,\tilde{{\bm{h}}}_{m,N+N_{p}}$ are not utilized downstream and their computation can therefore be bypassed.
@@ 2.2 Mixture of Experts
The Mixture of Experts (MoE) is a class of statistical machine learning frameworks that combines multiple models, known as experts, to produce more expressive and accurate predictions Jacob_Jordan-1991 ; jordan1994hierarchical . An MoE model typically consists of $N^{\prime}$ expert functions, $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{v}}$ for $i\in[N^{\prime}]$, along with a gating function, $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N^{\prime}}$ which assigns weights to the experts based on learned score functions, $s_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. For a given input ${\bm{h}}\in\mathbb{R}^{d}$, the MoE model generates its output as:
$\mathbf{y}=\sum_{j=1}^{N^{\prime}}G({\bm{h}})_{j}\cdot f_{j}({\bm{h}})=\sum_{j=1}^{N^{\prime}}\frac{\exp\left(s_{j}({\bm{h}})\right)}{\sum_{\ell=1}^{N^{\prime}}\exp\left(s_{\ell}({\bm{h}})\right)}\cdot f_{j}({\bm{h}}),$
where $G({\bm{h}})=\mathrm{softmax}(s_{1}({\bm{h}}),\ldots,s_{N^{\prime}}({\bm{h}}))$. Recent studies reveal connections between the attention mechanism, prompt-based methods, and MoE frameworks le2024mixture ; le2024revisiting , presenting new promising opportunities to investigate prompt-based techniques through MoE lens.
@@ 3 Motivation
Mixture of Experts meets Visual Prompt Tuning. We begin by discussing the connection between VPT and MoE. Adopting the notation from Section 2.1 and referring to Equation (2), let ${\bm{X}}^{(l)}=\left[{{\bm{x}}_{1}^{(l)}}^{\top},\dots,{{\bm{x}}_{N}^{(l)}}^{\top}\right]^{\top}\in\mathbb{R}^{Nd}$ be the concatenation of all $N$ input tokens to the $l$-th MSA layer. For notational simplicity, the layer superscript $(l)$ is omitted in the subsequent derivations. As shown in recent work le2024mixture , each output vector $\tilde{{\bm{h}}}_{m,i}$ within the $m$-th attention head can be equivalently expressed as the output of an MoE model with input ${\bm{X}}$. Specifically, we define the set of experts as:
@@
for $i\in[N]$. Here, $E_{j}\in\mathbb{R}^{d\times Nd}$ is a two-dimensional matrix such that $E_{j}{\bm{X}}={\bm{x}}_{j}$. Finally, the output of the $m$-th attention head for the $i$-th token can be expressed as follows:
@@
These formulations reveal that each attention head in ViT implicitly encodes multiple MoE models, $\tilde{{\bm{h}}}_{m,i}$ for $i\in[N]$. This contrasts with conventional MoE layers Quoc-conf-2017 , where experts and their gating functions typically operate on individual token embeddings ${\bm{x}}_{i}$. In our formulation, the expert networks and their score functions process the entire input sequence ${\bm{X}}$. Furthermore, these initial experts $f_{1},\dots,f_{N}$ and score functions are inherently part of the pre-trained ViT, with their parameters embedded within the model weights (see Equations (3) and (5)), thereby constituting pre-trained experts. Meanwhile, the new prompt experts $f_{N+1},\dots,f_{N+N_{p}}$, whose learnable parameters are contained within prompts (see Equations (4) and (6)), are introduced to efficiently adapt the model to downstream tasks. Consequently, VPT can be viewed as an efficient method for fine-tuning these implicit MoE models by adding new prompt experts. These added experts effectively act as downstream task experts, enabling specialization for new tasks without retraining the entire model.
@@
Restricted Functional Expressiveness of Prompt Experts. Equation (4) reveals a key limitation of prompt experts $f_{N+1},\dots,f_{N+N_{p}}$. Although they are formally functions of ${\bm{X}}$, they are represented by fixed prompt vectors ${\bm{p}}_{1},\dots,{\bm{p}}_{N_{p}}$ that remain constant regardless of the input. While their associated score functions $s_{i,N+j^{\prime}}$ in Equation (6) are input-dependent linear functions, the functional form realized by the prompt expert itself is static. This departs from typical MoE usage in the literature, where experts are adaptive functions of the input. Furthermore, it contrasts with the pre-trained experts $f_{1},\dots,f_{N}$, which are linear functions of ${\bm{X}}$ (see Equation (3)), making them comparatively more expressive. We hypothesize that this limited flexibility may constrain the effectiveness of VPT as a fine-tuning strategy. Supporting this, prior work petrov2023prompting demonstrates that prompt tuning can only add a bias term to the output of an attention block, thereby restricting representational capacity. This observation naturally motivates the central question: Can the performance of visual prompt tuning be improved by making prompt experts more expressive? Addressing this, we propose a novel prompt formulation to enhance their expressiveness in Section 4.
@@
Balancing Expressivity and Efficiency. Despite the limitations noted above, one of the main advantages of the current prompt design in VPT is its simplicity. As indicated by Equation (4), each prompt expert requires only $d$ parameters, making it highly parameter-efficient. One might consider a naive linear design for a prompt expert $f_{N+j^{\prime}}:\mathbb{R}^{Nd}\rightarrow\mathbb{R}^{d_{v}}$, where $f_{N+j^{\prime}}({\bm{X}})=W_{j^{\prime}}^{\top}{\bm{X}}$. However, this approach would necessitate up to $Nd\times d_{v}$ parameters, drastically increasing storage and computational overhead compared to the current approach, especially given the high dimensionality $Nd$ of the input ${\bm{X}}$. This illustrates the core design challenge: enhancing prompt expert expressiveness without sacrificing the parameter efficiency that makes VPT attractive. Our objective is therefore to devise a prompt mechanism that effectively balances these competing requirements.